# Pessimism Principle Can Be Effective: Towards a Framework for Zero-Shot Transfer Reinforcement Learning

### v1

@@ 7 Additional Discussion
Secondly, our method is not restricted to tabular setting, and can be integrated with function approximation or policy gradient. For example, when combine function approximation for robust RL [60] or robust policy gradient [61], the global step becomes parameter aggregation as in [42].
@@ 8 Experimental Results
In this section, we aim to verify the effectiveness of our pessimism principle, and its ability to avoid negative transfer. More details and additional experiments results are provided in Section B.
@@ 8.1 Effectiveness of Pessimism Principle for Transfer Learning
We develop our simulation on the recycling robot problem [62, 44]. In this problem, a mobile robot powered by a rechargeable battery is tasked with collecting empty soda cans. The robot operates with two battery levels: low and high. It has two possible actions: (1) search for empty cans; (2) remain stationary and wait for someone to bring it a can. When the robot’s battery is low (high), it has a probability of $\alpha$($\beta$) of finding an empty can and maintaining its current battery level, receiving some high reward. If the robot searches but does not find any, it will deplete its battery completely and receive a large penalty. If the robot chooses to wait, it will remain at the same battery level and receive a relatively small reward.
@@
To illustrate the robustness and generalizability of our methods against target model uncertainty, we evaluate the robust value functions of the learned policies against an uncertainty set centered at the target domain, which represents the worst-case performance under the model uncertainty. Figure 1-right shows that both of our pessimistic methods outperform the non-robust baseline, indicating that our approach can maintain a high performance, even with uncertainty or perturbations in the target domain. We also conduct an ablation study to compare the robust value functions under different target domain uncertainty levels in Appendix B.1, to further illustrate the enhanced robustness of our approaches.
@@ Appendix A Additional Related Works
Contextual Reinforcement Learning: In contextual reinforcement learning (cRL), the transition dynamics are assumed to be governed by an underlying context [63, 64]. This context can represent physical properties—such as wind conditions [65] or the length of a pole in a balancing task [66, 67, 68]—or more abstract features that characterize environment dynamics [69]. [70] identify cRL as particularly relevant for studying zero-shot generalization in RL agents. The cRL framework allows for a systematic and principled analysis of how agents adapt to environmental changes by explicitly defining inter- and extrapolation distributions. Following the evaluation protocol introduced by [70], [68] examined the generalization capabilities of various model-free RL agents on a benchmark that incorporates different physical properties as context information. Their approach assumed a naive utilization of context by directly concatenating it with observations. In contrast, [71] proposed a hypernetwork-based method to enable adaptable RL agents. However, these approaches still fail to address the fundamental challenges discussed earlier.
Robust RL: Robust RL [15, 14] or DRO formulation [72] offer a pessimism principle and a robust framework for addressing environmental uncertainty, which optimizes the worst-case performance within an uncertainty set of environments and can offer a theoretically sound lower bound on target domain performance when the target domain lies within this set. However, its application in transfer learning has been under explored, as robust RL often exhibits excessive pessimism, resulting in overly conservative policies for the target domain. Recent advances have extensively studied robust RL as a standalone problem [73, 61, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, wang2024unified, wang2024robust, zhang2025model], but its integration into transfer learning remains largely untapped. Leveraging robust RL principles in transfer learning offers an opportunity to develop methods that are not only theoretically grounded but also resilient to uncertainties in target domains. This proposal seeks to explore research directions that harness the strengths of robust RL to create more effective, efficient, and reliable transfer learning frameworks.
Federated Learning
@@
FL is a decentralized RL approach that aims to optimize the overall performance of K agents [87]. In this setting, each agent performs local updates based on its own environment and periodically communicates with a central server to aggregate the local models, without sharing raw trajectories with each other. While FL has been extensively studied [42, 48, 49, 88], existing work focuses on optimizing the average performance across local environments, rather than addressing the more challenging problem of multi-domain transfer. A fundamental challenge in applying FL to transfer learning lies in the nature of update rules: standard FL methods rely on linear updates (non-robust operators), whereas our proposed methods involve non-linear robust updates account for uncertainties.
@@ B.1 Ablation Study: experiments on different target domain uncertainty levels
To further evaluate the robustness of our methods, we conduct an ablation study analyzing the impact of different levels of uncertainty. In this experiment, we systematically vary the test uncertainty set radius $R_{\text{test}}$ and measure the optimal performance in the worst case over these uncertainty sets for the learned policies. Four different methods (MDTL-Avg, MDTL-Max, Non-robust DR, and Non-robust MDTL-Max) are evaluated across five levels of uncertainty, defined by the uncertainty set radius $\{0.01,0.03,0.05,0.07,0.1\}$. All testing experiments are conducted for 3 times. We also provide an optimal policy as a benchmark, which is trained on the target domain using non-robust vanilla value iteration [62]. The means and standard deviations of robust values of each method are reported over training time steps.
@@ B.4 Experiments on Biased Aggregation
The unbiased assumption made in MDTL-Max is to facilitate the convergence analysis. Nevertheless, our convergence results can still be extended to the case that the existing bias can be controlled, e.g., the bias introduced by the max aggregation can be controlled through techniques like threshold-MLMC [58]. On the other hand, even if there is bias, as long as the expected proxy is still pessimistic, our transfer learning framework and pessimism guarantees still hold. This indicates that our methods are robust to the bias.
@@ B.6 Comparison to Related Robust RL Approaches
Note that our methods are based on distributionally robust RL, where our uncertainty set is constructed to account for the potential distributional shift. We thus compare our proposed methods with other related robust RL approaches. We numerically verify that our methods enjoy both convergence and performance guarantees, and present the results under the recycling robot environment, where we use adversarial action-robust RL [89] and distributionally robust RL [81] as baselines.
@@ Appendix G Model-Free Algorithm Design and Analysis
We then develop studies for model-free setting. We will focus on $l_{p}$-norm defined uncertainty set [45, 90]:
@@
with $q=\frac{1}{1-\frac{1}{p}}$. For popular choices of $p$, the optimization problem in (117) has a closed-form solution, specified in Table 5 [45].
@@
(Theorem 1 in [45]) Let $\mathcal{P}_{s,a}$ be the uncertainty set defined using the $l_{p}$-norm. For any vector $V$, the following relationship holds:
@@
where the last inequality is from the fact that $\kappa$ is $1$-Lipschitz [45].

### v2

@@ 7 Additional Discussion
Secondly, our method is not restricted to tabular setting, and can be integrated with function approximation or policy gradient. For example, when combine function approximation for robust RL (Zhou et al., 2023) or robust policy gradient (Wang & Zou, 2022), the global step becomes parameter aggregation as in (Jin et al., 2022).
@@ 8 Experimental Results
In this section, we aim to verify the effectiveness of our pessimism principle, and its ability to avoid negative transfer. More details and additional experiments results are provided in Appendix B.
@@ 8.1 Effectiveness of Pessimism Principle for Transfer Learning
We develop our simulation on the recycling robot problem (Sutton & Barto, 2018; Wang et al., 2023b). In this problem, a mobile robot powered by a rechargeable battery is tasked with collecting empty soda cans. The robot operates with two battery levels: low and high. It has two possible actions: (1) search for empty cans; (2) remain stationary and wait for someone to bring it a can. When the robot’s battery is low (high), it has a probability of $\alpha$($\beta$) of finding an empty can and maintaining its current battery level, receiving some high reward. If the robot searches but does not find any, it will deplete its battery completely and receive a large penalty. If the robot chooses to wait, it will remain at the same battery level and receive a relatively small reward.
@@
To illustrate the robustness and generalizability of our methods against target model uncertainty, we evaluate the robust value functions of the learned policies against an uncertainty set centered at the target domain, which represents the worst-case performance under the model uncertainty. Figure 1-right shows that both of our pessimistic methods outperform the non-robust baseline, indicating that our approach can maintain a high performance, even with uncertainty or perturbations in the target domain. We also conduct an ablation study to compare the robust value functions under different target domain uncertainty levels in Section B.1, to further illustrate the enhanced robustness of our approaches.
@@ 9 Conclusion
Acknowledgment
This work was supported by DARPA under Agreement No. HR0011-24-9-0427 and NSF under Award CCF-2106339. The authors thank the anonymous reviewers whose constructive comments led to substantial improvement to the paper.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Additional Related Works
Imitation Learning and Policy Distillation: Imitation learning (IL) and policy distillation are extensively explored TL techniques that enable an agent to learn from demonstrations or from simpler policies. In IL, an agent mimics expert behavior through supervised learning, often using behavioral cloning or inverse RL (Hua et al., 2021; Desai et al., 2020; Kim et al., 2023; Zare et al., 2024). Policy distillation transfers knowledge from a teacher policy to a student policy by matching the teacher’s action distribution (Rusu et al., 2015; Yim et al., 2017; Yin & Pan, 2017; Traoré et al., 2019b, a). The success of these approaches, however, rely on the assumption that the source domain’s expert or teacher policy is highly aligned with the target domain, which may not always be the case.
@@
Contextual Reinforcement Learning: In contextual reinforcement learning (cRL), the transition dynamics are assumed to be governed by an underlying context (Hallak et al., 2015; Modi et al., 2018). This context can represent physical properties—such as wind conditions (Koppejan & Whiteson, 2009) or the length of a pole in a balancing task (Seo et al., 2020; Kaddour et al., 2020; Benjamins et al., 2022)—or more abstract features that characterize environment dynamics (Biedenkapp et al., 2020). Kirk et al. (2023b) identify cRL as particularly relevant for studying zero-shot generalization in RL agents. The cRL framework allows for a systematic and principled analysis of how agents adapt to environmental changes by explicitly defining inter- and extrapolation distributions. Following the evaluation protocol introduced by Kirk et al. (2023b), Benjamins et al. (2022) examined the generalization capabilities of various model-free RL agents on a benchmark that incorporates different physical properties as context information. Their approach assumed a naive utilization of context by directly concatenating it with observations. In contrast, Beukman et al. (2024) proposed a hypernetwork-based method to enable adaptable RL agents. However, these approaches still fail to address the fundamental challenges discussed earlier.
@@
Federated Learning: FL is a decentralized RL approach that aims to optimize the overall performance of K agents (McMahan et al., 2017). In this setting, each agent performs local updates based on its own environment and periodically communicates with a central server to aggregate the local models, without sharing raw trajectories with each other. While FL has been extensively studied (Jin et al., 2022; Wang et al., 2024b; Woo et al., 2023; Khodadadian et al., 2022), existing work focuses on optimizing the average performance across local environments, rather than addressing the more challenging problem of multi-domain transfer. A fundamental challenge in applying FL to transfer learning lies in the nature of update rules: standard FL methods rely on linear updates (non-robust operators), whereas our proposed methods involve non-linear robust updates account for uncertainties.
Robust RL: Robust RL (Iyengar, 2005; Nilim & El Ghaoui, 2004) or DRO formulation (Wiesemann et al., 2014) offer a pessimism principle and a robust framework for addressing environmental uncertainty, which optimizes the worst-case performance within an uncertainty set of environments and can offer a theoretically sound lower bound on target domain performance when the target domain lies within this set. However, its application in transfer learning has been under explored, as robust RL often exhibits excessive pessimism, resulting in overly conservative policies for the target domain. Recent advances have extensively studied robust RL as a standalone problem (Wang & Zou, 2021, 2022; Wang et al., 2023a; Badrinath & Kalathil, 2021; Dong et al., 2022; Lu et al., 2024; Liu & Xu, 2024; Yang et al., 2021; Xu et al., 2023; Panaganti & Kalathil, 2022; Shi et al., 2023; Wang et al., 2024a; Panaganti et al., 2022; Yang et al., 2022; Liu et al., 2023; Wang et al., 2024c, d; Zhang et al., 2025), but its integration into transfer learning remains largely untapped. Leveraging robust RL principles in transfer learning offers an opportunity to develop methods that are not only theoretically grounded but also resilient to uncertainties in target domains. This proposal seeks to explore research directions that harness the strengths of robust RL to create more effective, efficient, and reliable transfer learning frameworks.
@@ B.1 Ablation Study: experiments on different target domain uncertainty levels
To further evaluate the robustness of our methods, we conduct an ablation study analyzing the impact of different levels of uncertainty. In this experiment, we systematically vary the test uncertainty set radius $R_{\text{test}}$ and measure the optimal performance in the worst case over these uncertainty sets for the learned policies. Four different methods (MDTL-Avg, MDTL-Max, Non-robust DR, and Non-robust MDTL-Max) are evaluated across five levels of uncertainty, defined by the uncertainty set radius $\{0.01,0.03,0.05,0.07,0.1\}$. All testing experiments are conducted for 3 times. We also provide an optimal policy as a benchmark, which is trained on the target domain using non-robust vanilla value iteration (Sutton & Barto, 2018). The means and standard deviations of robust values of each method are reported over training time steps.
@@ B.4 Experiments on Biased Aggregation
The unbiased assumption made in MDTL-Max is to facilitate the convergence analysis. Nevertheless, our convergence results can still be extended to the case that the existing bias can be controlled, e.g., the bias introduced by the max aggregation can be controlled through techniques like threshold-MLMC (Wang et al., 2024f). On the other hand, even if there is bias, as long as the expected proxy is still pessimistic, our transfer learning framework and pessimism guarantees still hold. This indicates that our methods are robust to the bias.
@@ B.6 Comparison to Related Robust RL Approaches
Note that our methods are based on distributionally robust RL, where our uncertainty set is constructed to account for the potential distributional shift. We thus compare our proposed methods with other related robust RL approaches. We numerically verify that our methods enjoy both convergence and performance guarantees, and present the results under the recycling robot environment, where we use adversarial action-robust RL (Tessler et al., 2019) and distributionally robust RL (Panaganti & Kalathil, 2022) as baselines.
@@ Appendix G Model-Free Algorithm Design and Analysis
We then develop studies for model-free setting. We will focus on $l_{p}$-norm defined uncertainty set (Kumar et al., 2023b; Derman et al., 2021):
@@
with $q=\frac{1}{1-\frac{1}{p}}$. For popular choices of $p$, the optimization problem in (117) has a closed-form solution, specified in Table 5 (Kumar et al., 2023b).
@@
(Theorem 1 in (Kumar et al., 2023b)) Let $\mathcal{P}_{s,a}$ be the uncertainty set defined using the $l_{p}$-norm. For any vector $V$, the following relationship holds:
@@
where the last inequality is from the fact that $\kappa$ is $1$-Lipschitz (Kumar et al., 2023b).